A chemist needs 130 milliliters of a 31% solution but has only 17% and 43% solutions availble. Find how many millimiters of each that should be mixed to get the desired solution.
Answer by Alan3354(69443)

A chemist needs 130 milliliters of a 31% solution but has only 17% and 43% solutions availble. Find how many millimiters of each that should be mixed to get the desired solution.
------------------------
f = amount of 43%
s = amount of 17%
---------------
f + s = 130 (total solution)
0.43f + 0.17s = 130*0.31 (total of what's in the solution)
-------------
43f + 17s = 4030
s = 130 - f
-------------
43f + 17(130-f) = 4030
26f = 1820
f = 70 ml
s = 60 ml